a soccer player kicks a ball at an angle of 43 degrees with a speed of 10 m per second how far from the initial position will the ball land
1
Expert's answer
2021-10-14T18:39:07-0400

Using the formula for range from here https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/ , obtain:


"R = \dfrac{u^2\sin2\theta}{g}"

where "u = 10m\/s" is the initial speed, "\theta=43\degree" is the launch angle, "g = 9.8m\/s^2" is the gravitational acceleration. Thus, have:


"R = \dfrac{(10m/s)^2\cdot \sin(2\cdot 43\degree)}{9.8m/s^2} \approx 10m"

Answer. 10m.